# On the Convergence of Tâtonnement for Linear Fisher Markets

### v1

@@ 4.3 Linear Convergence to Approximate Equilibrium
For any $p$ such that $\lVert p-p^{*}\rVert\leq\epsilon\min_{j}p^{*}_{j}$, there exists an allocation $x$ such that $(p,x)$ constitute an $\epsilon$-approximate equilibrium.
@@
$\lVert p-p^{*}\rVert\leq\epsilon\min_{j}p^{*}_{j}$ implies $(1-\epsilon)p^{*}_{j}\leq p_{j}\leq(1+\epsilon)p^{*}_{j}$ for all $j$. Let $x^{*}$ be any equilibrium allocation associated with $p^{*}$. For any $i$, let $x_{ij}=x^{*}_{ij}$ if $p_{j}\leq p^{*}_{j}$ and $x_{ij}=(1-\epsilon)x^{*}_{ij}$ otherwise. Then, one can check:
@@
    $\langle v_{i},x_{i}\rangle\geq(1-\epsilon)\langle v_{i},x^{*}_{i}\rangle\geq(1-\epsilon)\langle v_{i},y_{i}\rangle$ for all $y_{i}$ such that $\langle p,y_{i}\rangle\leq(1-\epsilon)B_{i}$;
@@
    $\sum_{j}p_{j}x_{ij}\leq\sum_{j:p_{j}\leq p^{*}_{j}}p^{*}_{j}x^{*}_{ij}+\sum_{j:p_{j}>p^{*}_{j}}(1+\epsilon)p^{*}_{j}\cdot(1-\epsilon)x^{*}_{ij}\leq\sum_{j}p^{*}_{j}x^{*}_{ij}=B_{i}$;
@@
    $11-\epsilon\leq\sum_{i}(1-\epsilon)x^{*}_{ij}\leq\sum_{i}x_{ij}\leq\sum_{i}x^{*}_{ij}=$ 1.
@@
Remark 1.
@@
Zamani and Glineur (2023) showed a nearly-optimal last-iterate convergence rate without regularity conditions. Assuming that the function $\varphi(p)$ has $G$-bounded subgradients (i.e. any subgradient $g$ satisfies $\lVert g\rVert\leq G$) and $\lVert p^{0}-p^{*}\rVert\leq R$, they show that a sequence of iterates $\{p^{t}\}_{t=0}^{T}$ generated by the (projected) subgradient method with a constant stepsize $\frac{1}{\sqrt{T}}$ satisfies
$\varphi(p^{T})-\varphi(p^{*})\leq\frac{GR}{\sqrt{T}}\left(\frac{5}{4}+\frac{1}{4}\log{T}\right).$ (22)
@@
Note that smaller constant stepsizes than $\frac{1}{\sqrt{T}}$ do not guarantee the above convergence. As discussed before, for any $p^{0}\in\mathbb{R}^{m}_{\geq 0}$ such that $\sum_{j=1}^{m}p^{0}_{j}=B$, we have $\lVert p^{0}-p^{*}\rVert\leq\lVert p^{0}-p^{*}\rVert_{1}\leq\lVert p^{0}\rVert_{1}+\lVert p^{*}\rVert_{1}=2B$. Also, by Corollary 1, we have $\lVert g^{t}\rVert\leq\frac{B}{\min_{j}\tilde{p}_{j}-2m\eta}+m\leq\frac{2B}{\min_{j}\tilde{p}_{j}}+m$ for any $g^{t}\in\partial\varphi(p^{t})$ for all $t\geq 0$ if $\eta\leq\frac{\min_{j}\tilde{p}_{j}}{4m}$. Let $G=\frac{2B}{\min_{j}\tilde{p}_{j}}+m$ and $R=2B$. By combining our results with the results in Zamani and Glineur (2023), we obtain that after $T=\frac{G^{2}R^{2}}{\epsilon^{2}}\log^{2}\left(\frac{GR}{\epsilon}\right)$ iterations, tâtonnement with stepsize $\eta=\frac{1}{\sqrt{T}}=\frac{\epsilon}{GR\log(\frac{GR}{\epsilon})}$ generates a price vector $p$ satisfying $\varphi(p)-\varphi(p^{*})\leq\epsilon$, if $\epsilon\leq GR\min\{e^{-4},\frac{\min_{j}\tilde{p}_{j}}{m}\}$. See Section D.1 for a proof for this time complexity result.
@@
Using the subgradient inequality and our subgradient bound, it is also possible to use Eq. 17 to directly derive the duality gap convergence rate obtained in Remark 1. Conversely, one can apply the quadratic growth condition and the rate in Remark 1 to get a rate of convergence on $\lVert p^{t}-p^{*}\rVert^{2}$, however, this yields a rate of $\mathcal{\tilde{O}}(\frac{1}{\sqrt{T}})$, which is worse than the $\mathcal{\tilde{O}}(\frac{1}{T})$ rate in Eq. 17. Moreover, it does not allow one to obtain the linear rate of convergence to a neighborhood of the equilibrium, as in Eq. 17.
4.4 A Non-Convergence Example and Discussion.
Eq. 17 showed linear convergence to a neighborhood of the equilibrium prices, whose size depends on the stepsize $\eta$. Moreover, Corollary 2 showed the complexity for getting arbitrarily close to the equilibrium prices, but it requires choosing stepsizes accordingly. A natural question is then whether the last-iterate prices, under a fixed stepsize, converge to the equilibrium prices metrically or under any norm, i.e., $p^{t}\rightarrow p^{*}$ or $\|p^{t}-p^{*}\|\rightarrow 0$. The non-convergence of tâtonnement under linear utilities has been discussed in previous work, e.g., Cole and Fleischer (2008, Section 5), Cheung et al. (2019, Section 6), Avigdor-Elgrabli et al. (2014, Section 6). Cole and Tao (2019) gives a simple market instance (Example 1 in that paper) showing that a multiplicative form of tâtonnement results in prices cycling between off-equilibrium prices. Here, we provide a similar, simple instance to demonstrate the non-convergence of the additive form of tâtonnement Eq. 3 under any stepsize $\eta>0$. Consider a single buyer with a unit budget and two items with valuations $v_{11}=v_{12}=0.5$. The initial prices are chosen as $p^{0}=(0.5,0.5+\eta)$. It can be easily verified that the equilibrium prices $p^{*}=(0.5,0.5)$. For a small $\eta>0$, tâtonnement gives
and, as can be verified via straightforward calculations, $p_{1}^{2t+1}\rightarrow 0.5$, $p_{1}^{2t}\rightarrow 0.5+\eta$.
5 Tâtonnement in Quasi-linear Fisher markets
As a natural extension of the linear Fisher market, the quasi-linear Fisher market (QLFM) considers buyers who have outside value for money. Formally, given a set of prices $p=\{p_{j}\}_{j\in[m]}\geq 0$ on all items, the demand of buyer $i$ is
$\mathcal{D}^{q}_{i}(p)=\operatorname*{arg\,max}_{x_{i}}\left\{\langle v_{i}-p,x_{i}\rangle:\langle p,x_{i}\rangle\leq B_{i},\,x_{i}\geq 0\right\}.$
In the QLFM, if the price for every item is greater than the value to the buyer, the buyer prefers not to spend her money. Thus, unlike for LFM, buyers may not spend all their money in equilibrium. However, as in LFM, buyers only spend money on their MBB items, and the market clears in equilibrium.
As shown in Chen et al. (2007); Devanur et al. (2016), market equilibria in the QLFM can be captured by the following pair of EG-like primal and dual problems:
$\max_{x\in\mathbb{R}^{n\times m}_{\geq 0},u,y\in\mathbb{R}^{n}_{\geq 0}}\quad&\sum\nolimits_{i}\left(B_{i}\log{u_{i}}-y_{i}\right)\\ \textnormal{s.t.}\quad&u_{i}\leq\langle v_{i},x_{i}\rangle+y_{i}\quad\forall\,i\in[n]\\ &\sum\nolimits_{i}x_{ij}\leq 1\quad\forall\,j\in[m]\end{aligned}\quad\left|\quad\begin{aligned} \min_{\beta\in\mathbb{R}^{n}_{\geq 0},p\in\mathbb{R}^{m}_{\geq 0}}\quad&\sum\nolimits_{j}p_{j}-\sum\nolimits_{i}B_{i}\log{\beta_{i}}\\ \textnormal{s.t.}\quad&v_{ij}\beta_{i}\leq p_{j}\quad\forall\,i\in[n],j\in[m]\\ &\beta_{i}\leq 1\quad\forall\,i\in[n].\end{aligned}\right.$
Equivalence between tâtonnement and SubGD.
Similar to Eq. 1, the dual problem can be expressed as minimizing a max-type nonsmooth convex function
$\varphi^{q}(p)=\max_{x,y\geq 0}\sum_{i=1}^{n}\left(B_{i}\log{\left(\langle v_{i},x_{i}\rangle+y_{i}\right)}-y_{i}-\sum_{j=1}^{m}p_{j}x_{ij}\right)+\sum_{j=1}^{m}p_{j}.$ (23)
By the first-order optimality conditions, we have that for any $p$, any maximizer $\bar{x}$ of the inner maximization problem corresponds to a set of demands under price $p$, and the corresponding maximizer $\bar{y}$ denotes leftover budgets under $p$. Since any subgradient of $\varphi^{q}(p)$ is given by $1-\sum_{i=1}^{n}\bar{x}_{i}$, the subgradient equals the excess supply under price $p$.
By KKT conditions, the EG dual problem for QLFM can be simplified as
Again, the two expressions for $\varphi^{q}(p)$ only differ by a constant. Thus, we have $\partial\varphi^{q}(p)=-z^{q}(p)$ for any $p$. Consequently, tâtonnement in the QLFM is equivalent to SubGD on Eq. 24. See Appendix A for a direct derivation of this equivalence in terms of $\varphi^{q}(p)$.
Convergence analysis.
First, we show that the tâtonnement process in the QLFM generates a sequence of prices with lower and upper bounds, as long as the stepsize is not too large. Lower bounds on prices in the QLFM can be obtained almost identically to how we prove Lemma 1 since QLFM behaves the same as LFM when the price is small. Before showing the lemma, we introduce some notation. Recall that, given an item $j$, we use $\boldsymbol{b}(j)$ to denote the buyer with the minimum index such that
$\boldsymbol{b}(j)\in\operatorname*{\textnormal{Argmax}}_{i}\frac{v_{ij}}{\|v_{i}\|_{\infty}}.$
We denote $v_{\min}=\min_{j}v_{\boldsymbol{b}(j)j}$.
Lemma 5 (Lower bounds of prices in the QLFM).
Let $\hat{p}$ be an $m$-dimensional price vector where
Assume that we adjust prices in the QLFM with Eq. 3, starting from any initial price $p^{0}\geq\hat{p}$, with any stepsize $\eta<\frac{1}{2m}\min_{j}\hat{p}_{j}$. Then, $p^{t}_{j}\geq\hat{p}_{j}-2m\eta$ for all $j\in[m]$ and $t\geq 0$.
@@ 5 Tâtonnement in Quasi-linear Fisher markets
As in the proof of Lemma 1, we can show two facts analogous to Facts 1 and 10. Moreover, $\hat{p}_{j}<v_{\min}\max_{i}\frac{v_{ij}}{\lVert v_{i}\rVert_{\infty}}\leq v_{\boldsymbol{b}(j)j}$ guarantees buyer $\boldsymbol{b}(j)$ wants to spend all her money if $p_{j}\leq\hat{p}_{j}$. Other parts of the proof are analogous to the proof of Lemma 1. ∎
The complete proof Lemma 5 and all other omitted proofs for this section is in Appendix D. Completely analogously to the LFM case, we now get the following results.
Corollary 3 (Upper bounds of demands in the QLFM).
Assume that we adjust prices in the QLFM using Eq. 3, with $p^{0}\geq\hat{p}$ and $\eta<\frac{1}{2m}\min_{j}\hat{p}_{j}$. Let $z(p^{t})$ be the excess demand vector at time step $t$. Then, $\lVert z(p^{t})\rVert\leq\frac{B}{\min_{j}\hat{p}_{j}-2m\eta}+m$.
Lemma 6 (Upper bounds of prices in the QLFM).
Assume that we adjust prices in the QLFM using Eq. 3, with $p^{0}\geq\hat{p}$ and $\eta<\frac{1}{2m}\min_{j}\hat{p}_{j}$, where $\hat{p}_{j}$ is defined in Lemma 5. Then, $p^{t}_{j}\leq\bar{p}_{j}:=\min\{\max_{i}v_{ij},B\}+\frac{\eta B}{\hat{p}_{j}-2m\eta}$ for all $j\in[m]$ and $t\geq 0$.
Lemma 7 (Quadratic growth of $\varphi^{q}(p)$).
The convex function $\varphi^{q}(p)$ satisfies quadratic growth condition with modulus $\alpha=\min_{j}\frac{p^{*}_{j}}{2\big{(}\min\{\max_{i}v_{ij},B\}+\frac{\eta B}{\hat{p}_{j}-2m\eta}\big{)}^{2}}$, where $p^{*}$ denotes the unique equilibrium price vector and $\hat{p}$ is defined in Lemma 5.
Proof.
Let $(p^{*},x^{*},y^{*})$ be a market equilibrium in the QLFM and
$h^{q}(p)=\sum_{j=1}^{m}p_{j}-\sum_{i=1}^{n}\sum_{j=1}^{m}p^{*}_{j}x^{*}_{ij}\log{\left(\frac{p_{j}}{v_{ij}}\right)}.$ (26)
By KKT conditions of the EG-like primal and dual problems for QLFM, we have $\langle p^{*},x^{*}_{i}\rangle+y^{*}_{i}=B_{i}$ and $y^{*}_{i}>0$ implies $\min_{k\in[m]}\frac{p_{k}}{v_{ik}}\geq 1$ for all $i\in[n]$. Thereby, we have
$\varphi^{q}(p)$ $=\sum_{j=1}^{m}p_{j}-\sum_{i=1}^{n}\left(\sum_{j=1}^{m}p^{*}_{j}x^{*}_{ij}+y_{i}^{*}\right)\log{\left(\min\left\{\min_{k\in[m]}\frac{p_{k}}{v_{ik}},1\right\}\right)}$
$=\sum_{j=1}^{m}p_{j}-\sum_{i=1}^{n}\sum_{j=1}^{m}p^{*}_{j}x^{*}_{ij}\log{\left(\min\left\{\min_{k\in[m]}\frac{p_{k}}{v_{ik}},1\right\}\right)}-\sum_{i}y_{i}^{*}\log{\left(\min\left\{\min_{k\in[m]}\frac{p_{k}}{v_{ik}},1\right\}\right)}$
$=\sum_{j=1}^{m}p_{j}-\sum_{i=1}^{n}\sum_{j=1}^{m}p^{*}_{j}x^{*}_{ij}\log{\left(\min\left\{\min_{k\in[m]}\frac{p_{k}}{v_{ik}},1\right\}\right)}\geq h^{q}(p).$ (27)
Also, $\varphi^{q}(p^{*})=h^{q}(p^{*})$ since $x^{*}_{ij}>0$ implies $\frac{p^{*}_{j}}{v_{ij}}=\min_{k\in[m]}\frac{p^{*}_{k}}{v_{ik}}\leq 1$. Analogous to the LFM case, $\nabla_{p_{j}}h^{q}(p)=1-\frac{p^{*}_{j}}{p_{j}}\sum_{i=1}^{n}x^{*}_{ij}=1-\frac{p^{*}_{j}}{p_{j}}$ for all $j\in[m]$, thus $p=p^{*}$ is the minimizer of $h^{q}(p)$. Therefore, $\varphi^{q}(p)-\varphi^{q}(p^{*})\geq h^{q}(p)-h^{q}(p^{*})\geq\alpha\lVert p-p^{*}\rVert^{2}$, where $\alpha=\min_{j}\frac{p^{*}_{j}}{2\bar{p}_{j}^{2}}$. ∎
Based on the above results, we conclude a linear convergence rate and corresponding time complexity.
Theorem 3 (Linear convergence of tâtonnement in the QLFM).
$\lVert p^{t}-p^{*}\rVert^{2}\leq(1-2\eta\alpha)^{t}\lVert p^{0}-p^{*}\rVert^{2}+e\quad\text{for all }t\geq 0,$ (28)
where $\alpha$ is defined in Lemma 7 and
Corollary 4 (Time complexity of Tâtonnement in the QLFM).
For a given $\epsilon>0$, let $\eta\leq\min\left\{\frac{\min_{j}\hat{p}_{j}}{4m},\frac{2\epsilon\min_{j}p^{*}_{j}}{9B^{2}(\frac{2B}{\min_{j}\hat{p}_{j}}+m)^{2}}\right\}$. Then, starting from any $p^{0}\geq\hat{p}$ such that $\sum_{j}p^{0}_{j}=B$, tâtonnement with stepsize $\eta$ generates a price vector $p$ such that $\lVert p-p^{*}\rVert^{2}\leq\epsilon$ in $\mathcal{O}(\frac{1}{\epsilon}\log{\frac{1}{\epsilon}})$ iterations. If $\epsilon\geq\frac{9B^{2}\min_{j}\hat{p}_{j}}{8m\min_{j}p^{*}_{j}}(\frac{2B}{\min_{j}\hat{p}_{j}}+m)^{2}$, the time complexity is on the order of $\mathcal{O}(\log{\frac{1}{\epsilon}})$.
6 Numerical Experiments
In this section, we demonstrate the convergence of tâtonnement for various market settings through numerical experiments. We generate random market instances of various sizes for linear and quasilinear utilities and study the convergence of the tâtonnement prices toward equilibrium prices. To compute the “true” equilibrium prices, we run Proportional Response (PR) dynamics (Zhang, 2011) for LFM, and the quasi-linear variant of PR (Gao and Kroer, 2020), which gives highly accurate solutions for our simple instances.
We use the squared error norm, $\|p^{t}-p^{*}\|^{2}$, as the measure of convergence. For each market class, instance size, and step size, we run tâtonnement for a large number of iterations until the error residuals do not decrease further. In all instances, we run tâtonnement starting with $p^{0}_{j}=\frac{1}{m}$ for all $j\in[m]$, where $m$ is the number of items. We set the number of buyers and items to be $10,20,30,40$ and $20,40,60,80$, respectively.
Figure 1: Convergence of squared error norms on random generated instances of different sizes under linear utilities.
Figure 2: Convergence of squared error norms on random generated instances of different sizes under quasi-linear utilities.
Convergence to equilibrium prices under different stepsizes.
In Fig. 1 we plot the squared error norms $\|p^{t}-p^{*}\|^{2}$ across iterations for different LFM instance sizes. For each size, we test tâtonnement on a randomly generated instance: we draw budgets $\{B_{i}\}_{i\in[n]}$ from the $[0,1]$ uniform distribution and valuations $\{v_{i,j}\}_{i\in[n],j\in[m]}$ from the exponential distibution with scale $1$, and then normalize them to $\lVert B\rVert_{1}=1$ and $\lVert v_{i}\rVert_{1}=1$ for each $i$. Note that the y-axes are log-scaled. We compare the convergence of tâtonnement with three stepsizes: $2\times 10^{-5},4\times 10^{-5},6\times 10^{-5}$. The transparent dotted lines represent the values of $\|p^{t}-p^{*}\|^{2}$ for all iterations and the solid lines represent those values for equally sampled iterations. The corresponding plots for QL utilities are in Fig. 2.
Linear convergence.
For the same experiment settings, we zoom into the initial time steps of tâtonnement in Figs. 3 and 4. For each instance size, we only plot the line corresponding to stepsize $2\times 10^{-5}$ since "straight line" sections corresponding to different stepsizes are vastly different and don’t share time windows.
Figure 3: Linear convergence of squared error norms in the initial time steps of tâtonnement under linear utilities. All settings are the same as the corresponding plots in Fig. 1.
Figure 4: Linear convergence of squared error norms in the initial time steps of tâtonnement under quasi-linear utilities. All settings are the same as the corresponding plots in Fig. 2.
Discussion.
As can be seen from Figs. 1 and 2, under both linear utilities and quasi-linear utilities, for each instance size, tâtonnement with all three stepsizes ($2\times 10^{-5},4\times 10^{-5},6\times 10^{-5}$) lead to convergence of tâtonnement to small neighborhoods of the market equilibrium. Moreover, larger stepsizes lead to faster initial convergence rates but higher final error levels. This is consistent with our theoretical results: when $\eta\ll\min_{j}\tilde{p}_{j}-2m\eta$, then $\alpha\approx\min_{j}\frac{p^{*}_{j}}{2B^{2}}$ is approximately a constant, thus $1-2\eta\alpha$ linearly decreases as $\eta$ increases; and the final error level $e\approx\frac{B^{2}}{\min_{j}p^{*}_{j}}\eta\left(\frac{B}{\min_{j}\tilde{p}_{j}-2m\eta}+m\right)^{2}$ increases as $\eta$ increases.
As can be seen from Figs. 3 and 4, before oscillating around the final error levels, the lines corresponding to stepsize $2\times 10^{-5}$ exhibit an approximately linear convergence pattern. This demonstrates that our linear convergence results align with the practical performance of tâtonnement. That is to say, by considering the quadratic growth property of the convex potential function $\varphi(p)$, we discover the “true” underlying convergence rate of tâtonnement prior to hitting the oscillating neighborhood.
Acknowledgements
This research was supported by the Office of Naval Research awards N00014-22-1-2530 and N00014-23-1-2374, and the National Science Foundation awards IIS-2147361 and IIS-2238960.
Appendix A Direct equivalence between Tâtonnement and SubGD on CPF
Lemma 8.
Let
$f(x)=\min_{j}\mathbf{a}_{j}^{\top}x+b_{j}.$
Then, the subdifferential of $f(x)$ is
$\partial f(x)=\textnormal{conv}\left(\left\{\mathbf{a}_{j^{*}}\left|\,\mathbf{a}_{j^{*}}^{\top}x+b_{j^{*}}=\min_{j}\mathbf{a}_{j}^{\top}x+b_{j}\right.\right\}\right).$ (29)
Linear Fisher market.
Denote $\frac{1}{\mathbf{v}_{ij}}$ is an $m$-dimensional vector whose $j$-th entry is $\frac{1}{v_{ij}}$ and all other entries are 00. Let $\phi_{i}(p)=\min_{j\in[m]}\frac{p_{j}}{v_{ij}}=\min_{j\in[m]}\frac{1}{\mathbf{v}_{ij}}^{\top}p$. Denote $J^{*}_{i}(p)=\left\{j^{*}\left|\,\frac{p_{j^{*}}}{v_{ij^{*}}}=\min_{j\in[m]}\frac{p_{j}}{v_{ij}}\right.\right\}$. It follows that
$\partial\phi_{i}(p)=\text{conv}\left(\left\{\frac{1}{\mathbf{v}_{ij}}\Bigg{|}\,j\in J^{*}_{i}(p)\right\}\right)=\left\{\left(\frac{\lambda_{i1}}{v_{i1}},\ldots,\frac{\lambda_{im}}{v_{im}}\right)\left|\,\sum_{j=1}^{m}\lambda_{ij}=1,\lambda_{ij}\geq 0,\lambda_{ij}=0\text{ if }j\notin J^{*}_{i}(p)\right.\right\}.$
$=\mathbf{1}-\sum_{i}\left\{\frac{B_{i}}{\phi_{i}(p)}\left(\frac{\lambda_{1}}{v_{i1}},\ldots,\frac{\lambda_{m}}{v_{im}}\right)\left|\,\sum_{j=1}^{m}\lambda_{j}=1,\lambda_{j}\geq 0,\lambda_{j}=0\text{ if }j\notin J^{*}_{i}(p)\right.\right\}.$ (30)
Note that $B_{i}\lambda_{ij}=b_{ij}$ and $\phi_{i}(p)v_{ij}=p_{j}$ for $j\in J^{*}_{i}(p)$. Thus, $\partial\varphi(p)$ is equivalent to the excess supply (negative excess demand).
Quasi-linear Fisher market.
Let $\mathbf{a}_{i0}$ be an $m$-dimensional zero vector and $\mathbf{a}_{ij}=\frac{1}{\mathbf{v_{ij}}}$ for $j\in[m]$. Let $b_{i0}=1$ if $b_{ij}=0$ for $j\in[m]$. Let $\phi^{q}_{i}(p)=\min\left\{\min_{j\in[m]}\frac{p_{j}}{v_{ij}},1\right\}=\min_{j=0,1,2,\ldots,m}\mathbf{a}_{ij}^{\top}p+b_{ij}$. Denote $J^{*}_{i}(p)=\left\{j^{*}\left|\,\mathbf{a}_{ij^{*}}^{\top}p+b_{ij^{*}}=\min_{j=0,1,2,\ldots,m}\mathbf{a}_{ij}^{\top}p+b_{ij}\right.\right\}$. Here, $0\in J^{*}_{i}(p)$ means “do not spend money” is an optimal choice. It follows that
$\partial\phi^{q}_{i}(p)=\text{conv}\left(\left\{\mathbf{a}_{ij}\big{|}\,j\in J^{*}_{i}(p)\right\}\right)=\left\{\left(\frac{\lambda_{i1}}{v_{i1}},\ldots,\frac{\lambda_{im}}{v_{im}}\right)\left|\,\sum_{j=1}^{m}\lambda_{ij}\leq 1,\lambda_{ij}\geq 0;\lambda_{ij}=0\text{ if }j\notin J^{*}_{i}(p)\right.\right\}.$
$=\mathbf{1}-\sum_{i}\left\{\frac{B_{i}}{\phi^{q}_{i}(p)}\left(\frac{\lambda_{1}}{v_{i1}},\ldots,\frac{\lambda_{m}}{v_{im}}\right)\left|\,\sum_{j=1}^{m}\lambda_{j}\leq 1,\lambda_{j}\geq 0,\lambda_{j}=0\text{ if }j\notin J^{*}_{i}(p)\right.\right\}.$ (31)
Note that $B_{i}\lambda_{ij}=b_{ij}$ and $\phi^{q}_{i}(p)v_{ij}=p_{j}$ for $j\in[m]\cap J^{*}_{i}(p)$. Thus, $\partial\varphi^{q}(p)$ is equivalent to the excess supply (negative excess demand).
Appendix B Strong convexity of $h$
In this section, we compute the Hessian matrix of $h(p)$ and then determine the strong convexity modulus of $h(p)$. First, note that
$h(p)$ $=\sum_{j}p_{j}-\sum_{j}\sum_{i}p^{*}_{j}x^{*}_{ij}\log{\left(\frac{p_{j}}{v_{ij}}\right)}$
$=\sum_{j}p_{j}-\sum_{j}p^{*}_{j}\left(\sum_{i}x^{*}_{ij}\log{p_{j}}-\sum_{i}x^{*}_{ij}\log{v_{ij}}\right)$
$=\sum_{j}p_{j}-\sum_{j}p^{*}_{j}\left(\log{p_{j}}-\sum_{i}x^{*}_{ij}\log{v_{ij}}\right)$
$=\sum_{j}(p_{j}-p^{*}_{j}\log{p_{j}})-\sum_{i}\sum_{j}p^{*}_{j}x^{*}_{ij}\log{v_{ij}}.$
It follows that the Hessian matrix
$\nabla^{2}h(p)=\begin{bmatrix}\frac{p^{*}_{1}}{p_{1}^{2}}&0&\cdots&0\\ 0&\frac{p^{*}_{2}}{p_{2}^{2}}&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&\frac{p^{*}_{m}}{p_{m}^{2}}\end{bmatrix}\succeq\min_{j}\frac{p^{*}_{j}}{\bar{p}_{j}^{2}}\;\textbf{I},$ (32)
where I denotes the $m$-dimensional identity matrix. Equivalently, $h(p)$ is strongly convex with modulus $\mu=\min_{j}\frac{p^{*}_{j}}{\bar{p}_{j}^{2}}$. The strong convexity modulus of $h^{q}(p)$ can be computed in the same way.
Appendix C Algorithms
In this section, we show how to implement the discrete-time tâtonnement process in the linear Fisher market and the quasi-linear Fisher market. See Algorithms 1 and 2.
Input: initial price vector $p^{0}$
for $t=0,1,\ldots$ do
      $x^{t}_{j}\leftarrow 0$ for all $j$;
      for $i=1,2,\ldots,n$ do
           $J(i)=\operatorname*{\textnormal{Argmax}}_{j\in[m]}\frac{v_{ij}}{p^{t}_{j}}$;
           Set $\mbox{Set }b_{ij}\geq 0\;\forall\;j\in J(i)\mbox{ such that }\sum_{j\in J(i)}b_{ij}=B_{i}$;
           $x^{t}_{j}\leftarrow x^{t}_{j}+\frac{b_{ij}}{p^{t}_{j}}$ for $j\in J(i)$;
           $p^{t+1}_{j}=p^{t}_{j}+\eta(x^{t}_{j}-1)$ for all $j$;
ALGORITHM 1 Tâtonnement for the quasi-linear Fisher market
Input: initial price vector $p^{0}$
for $t=0,1,\ldots$ do
      $x^{t}_{j}\leftarrow 0$ for all $j$;
      for $i=1,2,\ldots,n$ do
           if $\max_{j\in[m]}\frac{v_{ij}}{p^{t}_{j}}>1$ then
                $J(i)=\operatorname*{\textnormal{Argmax}}_{j\in[m]}\frac{v_{ij}}{p^{t}_{j}}$;
                Set $\mbox{Set }b_{ij}\geq 0\;\forall\;j\in J(i)\mbox{ such that }\sum_{j\in J(i)}b_{ij}=B_{i}$;
                $x^{t}_{j}\leftarrow x^{t}_{j}+\frac{b_{ij}}{p^{t}_{j}}$ for $j\in J(i)$;
                $p^{t+1}_{j}=p^{t}_{j}+\eta(x^{t}_{j}-1)$ for all $j$;
ALGORITHM 2 Tâtonnement for the quasi-linear Fisher market
Appendix D Omitted proofs in Sections 4 and 5
D.1 Proof of the time complexity result in Remark 1.

### v2

@@ Appendix C More Detailed Discussion on the LFM
For any $p$ such that $\lVert p-p^{\textup{*}}\rVert\leq\epsilon\min_{j}p^{\textup{*}}_{j}$, there exists an allocation $x$ such that $(p,x)$ constitute an $\epsilon$-approximate equilibrium.
@@
$\lVert p-p^{\textup{*}}\rVert\leq\epsilon\min_{j}p^{\textup{*}}_{j}$ implies $(1-\epsilon)p^{\textup{*}}_{j}\leq p_{j}\leq(1+\epsilon)p^{\textup{*}}_{j}$ for all $j$. Let $*x^{\textup{*}}$ be any equilibrium allocation associated with $*p^{\textup{*}}$. For any $i$, let $x_{ij}=x^{\textup{*}}_{ij}$ if $p_{j}\leq p^{\textup{*}}_{j}$ and $x_{ij}=(1-\epsilon)x^{\textup{*}}_{ij}$ otherwise. Then, one can check:
@@
    $\langle v_{i},x_{i}\rangle\geq(1-\epsilon)\langle v_{i},x^{\textup{*}}_{i}\rangle\geq(1-\epsilon)\langle v_{i},y_{i}\rangle$ for all $y_{i}$ such that $\langle p,y_{i}\rangle\leq(1-\epsilon)B_{i}$;
@@
    $\sum_{j}p_{j}x_{ij}\leq\sum_{j:p_{j}\leq p^{\textup{*}}_{j}}p^{\textup{*}}_{j}x^{\textup{*}}_{ij}+\sum_{j:p_{j}>p^{\textup{*}}_{j}}(1+\epsilon)p^{\textup{*}}_{j}\cdot(1-\epsilon)x^{\textup{*}}_{ij}\leq\sum_{j}p^{\textup{*}}_{j}x^{\textup{*}}_{ij}=B_{i}$;
@@
    $11-\epsilon\leq\sum_{i}(1-\epsilon)x^{\textup{*}}_{ij}\leq\sum_{i}x_{ij}\leq\sum_{i}x^{\textup{*}}_{ij}=$ 1.
@@
Nearly-optimal last-iterate convergence rate without regularity conditions.
@@
Zamani and Glineur (2023) showed a nearly-optimal last-iterate convergence rate without regularity conditions. Assuming that the function $\varphi(p)$ has $G$-bounded subgradients (i.e. any subgradient $g$ satisfies $\lVert g\rVert\leq G$) and $\lVert p^{0}-p^{\textup{*}}\rVert\leq R$, they show that a sequence of iterates $\{p^{t}\}_{t=0}^{T}$ generated by the (projected) subgradient method with a constant stepsize $\frac{1}{\sqrt{T}}$ satisfies
$\varphi(p^{T})-\varphi(p^{\textup{*}})\leq\frac{GR}{\sqrt{T}}\left(\frac{5}{4}+\frac{1}{4}\log{T}\right).$ (26)
@@
Note that smaller constant stepsizes than $\frac{1}{\sqrt{T}}$ do not guarantee the above convergence. As discussed before, for any $p^{0}\in\mathbb{R}^{m}_{\geq 0}$ such that $\sum_{j=1}^{m}p^{0}_{j}=B$, we have $\lVert p^{0}-p^{\textup{*}}\rVert\leq\lVert p^{0}-p^{\textup{*}}\rVert_{1}\leq\lVert p^{0}\rVert_{1}+\lVert p^{\textup{*}}\rVert_{1}=2B$. Also, by Lemma 1 $(ii)$, we have $\lVert g^{t}\rVert\leq\frac{B}{\min_{j}\tilde{p}_{j}-2m\eta}+m\leq\frac{2B}{\min_{j}\tilde{p}_{j}}+m$ for any $g^{t}\in\partial\varphi(p^{t})$ for all $t\geq 0$ if $\eta\leq\frac{\min_{j}\tilde{p}_{j}}{4m}$. Let $G=\frac{2B}{\min_{j}\tilde{p}_{j}}+m$ and $R=2B$. By combining our results with the results in Zamani and Glineur (2023), we obtain that after $T=\frac{G^{2}R^{2}}{\epsilon^{2}}\log^{2}\left(\frac{GR}{\epsilon}\right)$ iterations, tâtonnement with stepsize $\eta=\frac{1}{\sqrt{T}}=\frac{\epsilon}{GR\log(\frac{GR}{\epsilon})}$ generates a price vector $p$ satisfying
$\varphi(p)-\varphi(p^{\textup{*}})\leq\epsilon,$ (27)
if $\epsilon\leq GR\min\{e^{-4},\frac{\min_{j}\tilde{p}_{j}}{m}\}$.
@@
Using the subgradient inequality and our subgradient bound, it is also possible to use Theorem 1 to directly derive the duality gap convergence rate obtained in Eq. 27. Conversely, one can apply the quadratic growth condition and the rate in Eq. 27 to get a rate of convergence on $\lVert p^{t}-p^{\textup{*}}\rVert^{2}$, however, this yields a rate of $\mathcal{\tilde{O}}(\frac{1}{\sqrt{T}})$, which is worse than the $\mathcal{\tilde{O}}(\frac{1}{T})$ rate in Theorem 1. Moreover, it does not allow one to obtain the linear rate of convergence to a neighborhood of the equilibrium, as in Theorem 1.
We show a proof for the above time complexity result.